OpenStudy (anonymous):

Martin and Anna buy books at a sale. Martin buys 3 hardcover books and 4 paperback books for $6.50. Anna buys 2 hardcover books and 6 paperback books for $6.00. What is the cost, in dollars, of each hardcover book?

OpenStudy (taylor<3srin):

Just think about it, it isn't hard.

OpenStudy (anonymous):

if the price of a hardcover is \(h\) and the price o a paperback is \(p\) then when we say Martin pays $6.50 for 3 hardcovers and 4 paperbacks this translates to \(3h+4p=6.5\) Similarly we say for Anna that \(2h+6p=6\). Can you solve this system of equations?$$3h+4p=6.5\\2h+6p=6$$

OpenStudy (anonymous):

is that the answer?